A chip in a microcomputer contain 5.72 x 10^-³ of silicon. Calculate the number of silicon atom in the ship (si=28; Avogadro constant = 6.02 × 10²³)
1
Expert's answer
2019-09-16T03:54:47-0400

If the mass is 5.72 * 10-3 grams, and the molar mass of silicon is 28, then the number of moles of silicon (5.72 * 10-3)/28 = 2,04*10-4. The Avogadro number shows the number of atoms in one mole. Then the number of atoms equals 2,04*10-4 * 6,02 * 1023 = 1,23*1020.